OpenStudy (anonymous):

Help. Giving Out Medals ;) PLEASE HELP!! A line passes through (-7,-5) and (-5,4). Write an equation for the line in point-slope form. Rewrite the equation in standard form using integers. A. y+5= 9/2 (x-7); -9x+2y=53 B. y+7= 9/2 (x+5); -9x+2y=31 C. y+5= 9/2 (x+7); -9x+2y=53 D. y-5= 9/2 (x+7); -9x + 2y=-53

OpenStudy (anonymous):



OpenStudy (anonymous):

m = y2 -y1/x2-x1 = (-5) - 4/(-7)-(-5) = -9/-2 = 9/2

OpenStudy (anonymous):

y - y1 =m ( x - x1) = 0 since y1 = -5 then: y - (-5) =9/2( x - (-5)) = 0 y + 5 = 9/2 (x + 5) sorry to tell you answers are wrong..it happens sometimes

OpenStudy (anonymous):

Need better help